solve 32-18m2(square)=0
Answer by checkley77(12844)

32-18m^2=0
2(16-9m^2)=0
2(4+3m)(4-3m)=0
4+3m=0\
3m=-4
m=-4/3 ans.
4-3m=0
-3m=-4
m=4/3 ans.